Consider the binary operation defined by x * y=xy-2x-15 Solve x *2=0
Found 2 solutions by josgarithmetic, ikleyn:
Answer by josgarithmetic(39617)

Substituting and following the definition

x*y=xy-2x-15
x*(2)=(2)(y)-2(2)-15=0
2y-4-15=0
2y=19
y=19/2------------------nineteen halves

Answer by ikleyn(52781)

.
Consider the binary operation defined by x * y = xy-2x-15. Solve x * 2=0
~~~~~~~~~~~~~~~~~~


        The solution in the post by @josgarithmetic is  WRONG.
        I came to bring a correct solution.


To get a tradition form of the x * 2 = equation,

we should substitute y= 2 in the expression  xy - 2x - 15.


By doing it, you get this equation

  2x - 2x - 15 = 0.


This equation is reduced to -15 = 0 after combining like term.


The last equation, -15 = 0, is self-contradictory and does not have solutions.


Therefore, the original equation x * 2 = 0 does not have solutions.


ANSWER.  The given equation does not have solution.

Solved.

-----------------

Ignore the post by @josgarithmetic, for the safety of your mind.